OpenStudy (anonymous):

Value of lim x->2 (1/(x-2)-4/(x^2)) indeterminate inf-inf so L'Hospital rule combine terms and I think I get 1-4/x-2-x^2-4?

OpenStudy (anonymous):

Difficult to understand the terms.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} (\frac{ 1 }{ x-2 }-\frac{ 4 }{ x^2-4 })\] combine to \[\lim_{x \rightarrow 2}(\frac{ 1-4 }{ x-2-x^2-4 })\]

OpenStudy (anonymous):

which I'm certain is wrong

OpenStudy (anonymous):

yeah 2 nd step is wrong.

OpenStudy (anonymous):

First write x^2-4 as (x+2)(x-2). Then take out 1/(x-2) out of the terms. You will be having\[1/(x-2) [ 1-4/(x+2)]\] now you can proceed.

OpenStudy (anonymous):

So I'm left with \[\lim_{x \rightarrow 2}(\frac{ 4 }{ x+2 })\]? That doesn't seem right or am I completely missing something?

OpenStudy (anonymous):

no, how are you still left with the four?

OpenStudy (anonymous):

oh so I've got -3/x+2?

OpenStudy (anonymous):

no, 1- 4/x+2 = x+2-4/x+2 =x-2/x+2 the x-2 gets cancelled. so all thats left is 1/x+2

OpenStudy (anonymous):

Aaaaaa- that helps. But I'm pretty sure that the question is asking for me to apply L'Hospital's Rule which I need to show the work on

OpenStudy (anonymous):

so the \[\lim_{x \rightarrow 2}=\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

L'hospital method can be applied only when you have a indeterminate form. The answer is correct.

OpenStudy (anonymous):

Thanks

OpenStudy (jhannybean):

\[\large \lim_{x \rightarrow 2} (\frac{ 1 }{ x-2 }-\frac{ 4 }{ x^2-4 })\]evaluate the inside.\[\large \frac{1}{x-2}-\cfrac{4}{(x-2)(x+2)}\]Make a common denominator. \[\large \cfrac{(x+2)-4}{(x+2)(x-2)}\]Simplify \[\large \cfrac{x-2}{(x+2)(x-2)}\]Simplify again. \[\large \cfrac{1}{x+2}\]Now take the limit of \(\cfrac{1}{x+2}\)\[\large \lim_{x \rightarrow 2}\cfrac{1}{x+2}\]\[\large =\cfrac{1}{4} \]